OpenStudy (anonymous):

3+(5n^2)/n+(n^2) converge or diverge

OpenStudy (agreene):

are you asking: \[\sum_{n=1}^{\infty}3+\frac{5n^2}{n}+n^2\] because thats how you have it written, and clearly it doesnt converge.

OpenStudy (anonymous):

no it 3+5n^2 over n+n^2

OpenStudy (anonymous):

its*

OpenStudy (agreene):

\[\sum_{n=1}^{\infty} \frac{3+5n^2}{n+n^2}\] diverges via limit test.

OpenStudy (anonymous):

and n to infinity not 1

OpenStudy (agreene):

that series is for n=1 to infinity. The n=1 subscript is the starting k and the top is the ending k.

OpenStudy (anonymous):

oh i see ... but the my book says that the final answer is 5 ?! :(

OpenStudy (agreene):

either I'm not understanding the question, or your book is wrong.

OpenStudy (anonymous):

mmm its ok thanks any way for the help :)